Table 2: Intellegence Rating of 100 students. Intellegence Rating Frequency 5 15 6 27 7 39...

  1. Table 2: Intellegence Rating of 100 students.

    Intellegence Rating

    Frequency

    5

    15

    6

    27

    7

    39

    8

    16

    9

    3

  2. If we assume the participants in the survey of 100 students are a “typical” sample of students. What is the probability of selecting a student with an intellegence rating of 6? Make sure to show your calculations for full credit.
  3. What is the probability of selecting student with a Intellegence Rating of 8 or 9? P(8 or 9). Use the Addition Rule. Show your work for full credit.
Answer #1

Answer:

11. 75% chances is the prediction of rain tomorrow,i.e.,P[raining tomorrow]=0.75

P[not raining tomorrow]=1-P[raining tomorrow]=1-0.75=0.25

12. P[selecting a student with rating 6]=,where n is the total number of students and is the frequency at 6.

Here,n=100

Hence,P[selecting a student with rating 6]=

13. P[selecting a student with rating of 8 or 9]=P[selecting a student with rating 8]+P[selecting a student with rating 9]

P[selecting a student with rating 8]=,where n is the total number of students and is the frequency at 8.

P[selecting a student with rating 9]=,where n is the total number of students and is the frequency at 9.

NOTE:

I HOPE THIS ANSWER IS HELPFUL TO YOU.

** PLEASE SUPPORT ME WITH RATING.

..... THANK YOU ....